# Photon propagation in a charged Bose-Einstein condensate

### v1

@@ 5 Discussion
$M^{2}=$ $=$ $\frac{m^{2}_{V}+m^{2}_{+}}{4\mu^{2}}\,,$
$\Delta=$ $=$ $\frac{|m^{2}_{V}-m^{2}_{+}|}{4\mu^{2}}\,.$ (5.1)
@@
Plots of the dispersion relations are shown in Fig. 1 for sample values of the mass parameters. It must be kept in mind that numerically $\Delta<M^{2}$, and also it is important to recognize that
@@
Eq. (5.3) follows from the formulas for $m^{2}_{V}$ and $m^{2}_{+}$ in terms of $\mu$ (e.g. Eqs. (4.16) and (4.17)), that give
@@
One particular feature worth mentioning is the result for the non-relativistic (NR) limit of the $\omega_{-}$ branch in the case that $\Delta\simeq 1$. For sufficiently small values of $\kappa$, the dispersion relations are
Thus, in particular,
$(y^{2}_{-})_{NR}=(M^{2}-\Delta)+\left(1-\Delta^{-1}\right)x^{2}+\frac{x^{4}}{2\Delta^{3}}\,,$ (5.6)
and in the case that $\Delta\simeq 1$, the dispersion of the ()(-)( - ) branch is
$(y^{2}_{-})_{NR}\simeq(M^{2}-1)+\frac{1}{2}x^{4}\,.$ (5.7)
In other words, the $\kappa^{2}$ does not appear, which is reminiscent of the dispersion relation that has been obtained in NR models of the BE condensation of charged scalars[21]. However, in our case, the (±)plus-or-minus($\pm$)( ± ) dispersion relations are different, and this occurs only for the $\omega_{-}$ dispersion relation. More generally, the presence of a $O(\kappa^{2})$ term (even in the $T=0$ regime that we are considering) is mentioned in Ref. [22], as well as the fact that the coefficient is negative; which in our case can occur for $\omega_{-}$ if $\Delta<1$.
Another noteworthy feature is the fact that if $\Delta<1$, then $\omega_{-}$ decreases in a certain range of $\kappa$, and reaching a minimum value. This is illustrated in Fig. 1 in the plots of $\omega_{-}$ for $\Delta=0.2,0.5$, and more emphatically in Fig. 2.
@@
In any case, the expression for $\omega^{2}_{-}$, as given by Eq. (5.2), remains positive for all values of $\kappa$. To show this let us note the following. From Eq. (5.2) it follows that $y^{2}_{-}$ becomes zero at $x=x_{\pm}$, where
@@
Plots of $v_{g}$ are shown in Figs. 3 and 4 for the $\omega_{\pm}$ branches, respectively. Each figure also displays the corresponding plot of $v_{\pm}\equiv\kappa/\omega_{\pm}$ for reference. In each branch, $v$ does not change appreciably for the different values of $\Delta$. The group velocity of the $\omega_{-}$ branch becomes negative in the range of $\kappa$ in which $\omega_{-}$ decreases. In fact it is zero at
@@
The results discussed in this section indicate that the BE condensate medium has non-trivial optical properties. For example, as we have observed above, the group velocity of the $\omega_{-}$ mode can be reduced, zero or negative, depending on the value of $\kappa$. In fact, it has been experimentally shown that the speed of light can be reduced drastically in a BE condensate[15], and the results can have consequences in other contexts. For example, if the dark matter has a BE condensate component, such situation may arise that due to the dispersive nature of the dark matter medium, the properties of the propagating photons may be drastically altered and maybe even that no electromagnetic signal may be observed within a certain range of $\kappa$.
@@ 6 Polarization tensor and dielectric constant
An alternative way to obtain the dispersion relations is through the photon polarization tensor, or self-energy. This method has also the virtue that it provides us with an expression for the dielectric function of the system, which is useful on its own merits. The starting point is the bilinear part of the Lagrangian for $V_{\mu}$ and $\rho$ given in Eq. (4.12). The idea is to integrate out the variable $\rho$, which is the tree-level analog of calculating the 1-loop self-energy, leaving only the dynamical variables of $V$, and identify the polarization tensor by comparing the result with
@@
which can be written in the form
$\pi_{L}=m^{2}_{V}+\frac{4\mu^{2}m^{2}_{V}\kappa^{2}}{k^{2}(k^{2}-m^{2}_{1})-4\mu^{2}(k\cdot u)^{2}}\,,$ (6.6)
remembering that
$\tilde{u}^{2}=-\frac{\kappa^{2}}{k^{2}}\,.$ (6.7)
@@ 7 Mixing
Besides the dispersion relation and the polarization tensor, another important factor is the $\rho-V_{L}$ mixing in the eigenmodes. Once the solutions $\omega_{\pm}$ are given, these are determined from Eq. (4.1).
@@
We write the elements of the mode eigenvectors in the form
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}$ $=$ $\frac{1}{\sqrt{N_{-}}}\left(\begin{array}[]{c}\Lambda\\[12.0pt] m_{V}\gamma_{-}\end{array}\right)\,,$ (7.5)
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{+}$ $=$ $\frac{1}{\sqrt{N_{+}}}\left(\begin{array}[]{c}-m_{V}\gamma_{+}\\[12.0pt] \Lambda-\gamma^{2}_{+}\end{array}\right)\,,$ (7.10)
@@
$k^{2}_{\pm}$ $=$ $\omega^{2}_{\pm}-\kappa^{2}=\frac{1}{2}(m^{2}_{V}+m^{2}_{+})\pm\left[\frac{1}{4}(m^{2}_{V}-m^{2}_{+})^{2}+4\mu^{2}\kappa^{2}\right]^{\frac{1}{2}}\,,$
$\Lambda$ $=$ $\frac{1}{2}(m^{2}_{V}-m^{2}_{+})+\left[\frac{1}{4}(m^{2}_{V}-m^{2}_{+})^{2}+4\mu^{2}\kappa^{2}\right]^{\frac{1}{2}}\,,$ (7.11)
@@
$N_{+}$ $=$ $(\Lambda-\gamma^{2}_{+})^{2}+(m_{V}\gamma_{+})^{2}\,,$
$N_{-}$ $=$ $\Lambda^{2}+(m_{V}\gamma_{-})^{2}\,.$ (7.12)
With this convention, when $m_{V}>m_{+}$ in the limit $\kappa=0$ these solutions correspond to
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}$ $=$ $\left(\begin{array}[]{c}1\\ 0\end{array}\right)\,,$ (7.17)
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{+}$ $=$ $\left(\begin{array}[]{c}0\\ 1\end{array}\right)\,.$ (7.22)
Plots of the mixing coefficients $|\rho_{\pm}|^{2}$ and $|V_{L\pm}|^{2}$ is shown in Fig. 5 and Fig. 6. For the plots we take $M^{2}$ and $\Delta$ as indicated in the caption, and we consider the case $m_{V}>m_{+}$ so that, in particular, $(m_{V}/2\mu)^{2}=\frac{1}{2}(M^{2}+\Delta)$. The plots do not vary significantly for other values of $M^{2}$ and $\Delta$.
As already indicated above, at $\kappa=0$ the $\omega_{-}$ solution corresponds to a pure $\rho$,
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}=\left(\begin{array}[]{c}1\\ 0\end{array}\right)\qquad(\kappa=0)\,.$ (7.23)
@@
$\kappa^{2}_{0}=\frac{m^{2}_{V}m^{2}_{+}}{4\mu^{2}}\,.$ (7.25)
It is easy to see from Eq. (7), that $\gamma_{-}$ increases relative to $\Lambda$ as $k^{2}_{-}\rightarrow 0$, and therefore from Eq. (7.5) it follows that
@@
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}=\frac{1}{\sqrt{N_{-}}}\left(\begin{array}[]{c}\Lambda\\[12.0pt] -im_{V}|\gamma_{-}|\end{array}\right)\qquad(\kappa>\kappa_{0})\,.$ (7.28)
@@
$e_{\pm}=\left(\begin{array}[]{c}a\\ \pm ib\end{array}\right)\,.$ (7.29)
@@ 8 Conclusions and Outlook
We have presented a model for the propagation of photons in a medium that contains a Bose-Einstein (BE) condensate of a charged scalar field. The idea is based on the model recently proposed to treat the propagation of fermions in a BE condensate, extended by coupling the scalar field with the electromagnetic field in the usual way. Our goal has been to study the effects of the symmetry breaking on the the photon dispersion relations and the optical properties of the system. To this end, the propagating modes were identified, and the corresponding dispersion relations were determined, in Section 4. As we have shown there, two of the modes correspond to the transverse photon polarizations, and their dispersion relations have the usual form of the transverse photons in a plasma. There are two other modes, which we denote as the (±)plus-or-minus($\pm$)( ± ) modes, which are combinations of the longitudinal photon and the massive scalar field. Some salient aspects of the results obtained were discussed in Section 5. A particularly distinguishable one is the fact that the dispersion relation of the ()(-)( - ) mode decreases as a function of the momentum in a given range, and therefore the corresponding group velocity is negative in that range, which shows that the BE condensate can have some peculiar and non-trivial properties. As a complement, in Section 6 we gave the expression for the photon polarization tensor. On one hand, this provides an alternative way to obtain the dispersion relations, but more importantly it has also the virtue that it provides an expression for the dielectric function of the system, which is useful on its own merits, e.g. for studying also the response of the system to external fields. Another useful result is the determination of the mixing parameters of the (±)plus-or-minus($\pm$)( ± ) modes presented in Section 7, which can be used to obtain the corrections to the dispersion relations (e.g., imaginary parts due the damping effects) and/or the effects of scattering, due to the interactions with the excitations of the system. In summary, the optical properties of a medium that contains a BE condensate have particular features that motivate further exploration. The results of the present work can be useful to study the effects they may have in specific applications in the various contexts considered in the references cited.
@@
What is the Lagrangian in momentum space? Using the transformation of $a$ and $b$,
@@
In particular that this $L(k)$ is not real (Hermitian). The correct procedure is to notice that, since the first term mixes $a(k)$ with $a(-k)$, and similarly with $b(k)$, we must include both also in the $\alpha$ term. Therefore we must write

### v2

@@ 5 Discussion
$M^{2}$ $=$ $\frac{m^{2}_{\rho}+m^{2}_{V}}{4\mu^{2}}\,,$
$\Delta$ $=$ $\frac{|m^{2}_{\rho}-m^{2}_{V}|}{4\mu^{2}}\,.$ (5.1)
@@
For guidance, plots of the dispersion relations are shown in Fig. 1 for sample values of the mass parameters. It must be kept in mind that numerically $\Delta<M^{2}$, and also it is important to recognize that
@@
Eq. (5.3) follows from the formulas for $m^{2}_{V}$ and $m^{2}_{\rho}$ in terms of $\mu$ (e.g. Eqs. (4.17) and (4.16)), that give
@@
From Eq. (5.2) it follows that $y_{+}$ increases steadily as $x$ increases, and $y_{+}>x$, for all values of $x$. In other words $\omega_{+}$ increases steadily with $\kappa$, and $\omega_{+}>\kappa$, for all values of $\kappa$. However, $\omega_{-}$ has some distinguishing features, and its behavior is different depending on the range of the value of $\kappa$. The plots in Figs. 2 and 3 show the dispersion relations covering different regimes of $\kappa$ to emphasize some of these features.
Firstly, it follows from Eq. (5.2) that
$y_{-}<x$ (5.5)
for
$x^{2}>\frac{1}{2}(M^{4}-\Delta^{2})\,.$ (5.6)
That is,
$\omega_{-}<\kappa\,,$ (5.7)
for
$\frac{\kappa^{2}}{\mu^{2}}>M^{4}-\Delta^{2}\,,$ (5.8)
as illustrated in Fig. 2. The dispersion relation in the $\kappa$ regime indicated in Eq. (5.8) is reminiscent of the photon dispersion relation in a medium with an index of refraction greater than 1.
Figure 2: Sample plots of the dispersion relations to illustrate that $\omega_{-}<\kappa$ when $\kappa$ is sufficiently large; i.e., Eqs. (5.6) and (5.8). For reference the line $y=x$ is shown as well. We have set $M=2$, and $\Delta$ as indicated.
@@
Secondly, another noteworthy feature is the fact that if $\Delta<1$, then $\omega_{-}$ decreases in a certain range of $\kappa$, and reaching a minimum value. This is illustrated in Fig. 1 in the plots of $\omega_{-}$ for $\Delta=0.2,0.5$, and more emphatically in Fig. 3. In any case, the expression for $\omega^{2}_{-}$, as given by Eq. (5.2), remains positive for all values of $\kappa$. To show this let us note the following. From Eq. (5.2) it follows that $y^{2}_{-}$ becomes zero at $x=x_{\pm}$, where
@@
Plots of $v_{g}$ are shown in Figs. 4 and 5 for the $\omega_{\pm}$ branches, respectively. Each figure also displays the corresponding plot of $v_{\pm}\equiv\kappa/\omega_{\pm}$ for reference. In each branch, $v$ does not change appreciably for the different values of $\Delta$. The group velocity of the $\omega_{-}$ branch becomes negative in the range of $\kappa$ in which $\omega_{-}$ decreases. In fact it is zero at
@@
One particular feature worth mentioning is the result for the non-relativistic (NR) limit of the $\omega_{-}$ branch. For sufficiently small values of $\kappa$, the dispersion relations are
Thus, in particular,
$(y^{2}_{-})_{NR}=(M^{2}-\Delta)+\left(1-\Delta^{-1}\right)x^{2}+\frac{x^{4}}{2\Delta^{3}}\,.$ (5.14)
Moreover, in the case that $\Delta\simeq 1$, the dispersion of the ()(-)( - ) branch is
$(y^{2}_{-})_{NR}\simeq(M^{2}-1)+\frac{1}{2}x^{4}\,.$ (5.15)
To state this more precisely, recall that in the NR limit, the chemical potential, which we denote by $\mu_{nr}$, is related to the chemical potential that we use in TFT by
$\mu=m+\mu_{nr}\,.$ (5.16)
This follows from the usual form of the distribution functions,
$f(p)=\frac{1}{e^{\beta(E(p)-\mu)}\pm 1}\,,$ (5.17)
and from the NR counterpart
$f_{NR}(p)=\frac{1}{e^{\beta(p^{2}/2m-\mu_{nr})}\pm 1}\,.$ (5.18)
Let us consider specifically the case
$m\gg\kappa,\mu_{nr},m_{V}\,.$ (5.19)
In that limit Eq. (5.14) gives
$(y^{2}_{-})_{NR}=\frac{m^{2}_{V}}{2\mu^{2}}+x^{2}O\left(\frac{\mu_{nr}}{m},\frac{m^{2}_{V}}{m^{2}}\right)+\frac{1}{2}x^{4}\left(1+O\left(\frac{\mu_{nr}}{m},\frac{m^{2}_{V}}{m^{2}}\right)\right)\,,$ (5.20)
or, in terms of the physical variables,
$(\omega^{2}_{-})_{NR}=m^{2}_{V}+\left(\frac{\kappa^{2}}{2m}\right)^{2}+O\left(\frac{\kappa^{2}\mu_{nr}}{m},\frac{\kappa^{2}m^{2}_{V}}{m^{2}}\right)\,.$ (5.21)
Thus if the $O\left(\kappa^{2}\mu_{nr}/m,\kappa^{2}m^{2}_{V}/m^{2}\right)$ terms can be neglected, the $\kappa^{2}$ term does not appear, which is reminiscent of the dispersion relation that has been obtained in NR models of the BE condensation of charged scalars[21]. However, in our case, the (±)plus-or-minus($\pm$)( ± ) dispersion relations are different, and this occurs only for the $\omega_{-}$ dispersion relation. More generally, the presence of a $O(\kappa^{2})$ term (even in the $T=0$ regime that we are considering) is mentioned in Ref. [22], as well as the fact that the coefficient is negative; which in our case can occur for $\omega_{-}$ if $\Delta<1$.
@@ 6 Polarization tensor and dielectric constant
An alternative way to obtain the dispersion relations is through the photon polarization tensor, or self-energy. This method has also the virtue that it provides us with an expression for the dielectric function of the system, which is useful on its own merits. The starting point is the bilinear part of the Lagrangian density for $V_{\mu}$ and $\rho$ given in Eq. (4.12). The idea is to integrate out the variable $\rho$, which is the tree-level analog of calculating the 1-loop self-energy, leaving only the dynamical variables of $V$, and identify the polarization tensor by comparing the result with
@@
which using Eqs. (2), (4.1) and (4.16) can be written in the form
$\pi_{L}=m^{2}_{V}+\frac{4\mu^{2}m^{2}_{V}\kappa^{2}}{k^{2}(k^{2}-m^{2}_{\rho})-4\mu^{2}\kappa^{2}}\,.$ (6.6)
@@
In this context we would like to emphasize the following. What we have done is to find the modes that diagonalize the scalar-photon part of the Lagrangian, in the BE background. The dispersion relations that we have written in Eq. (4.20) are the dispersion relations of those modes.
In reality, as we have mentioned, the BE condensate must be superimposed on an oppositely charged background component. It would be a multi-component system, of which one of them is the BE condensate. Both components contribute to the total $\pi_{L}$ (as well as the dielectric constant), and therefore the dispersion relations of the propagating plasma modes have contributions from both components. Physically the propagating modes are combinations of the BE modes and the plasma modes of the background component.
In systems in which the dominant contribution to $\pi_{L}$ arises from the BE component [which we have written in Eq. (6.6)], then the dispersion relation of the propagating photon (obtained by solving Eq. (6.7)) is given by the dispersion relation of the BE modes that we have written in Eq. (4.20).
In more general cases, $\pi_{L}$ would have an additional contribution besides the BE contribution of Eq. (6.6), depending on the composition and conditions of the background component, and the dispersion relations of the propagating modes would be correspondingly different from those given in Eq. (4.20).
The main result is that, in such a multicomponent system, Eq. (6.6) represents the contribution to $\pi_{L}$ (and whence to the dielectric constant) due to the BE component, which can be used to study the implications for the optical properties of the system, which are different from what they would be if the system does not have the BE component.
@@ 7 Mixing
Besides the dispersion relation and the polarization tensor, another important factor is the $\rho-V_{L}$ mixing in the eigenmodes. Once the solutions $\omega_{\pm}$ are given, these are determined from Eq. (4.1). From Eq. (4.20) for the dispersion relations it follows that for $\kappa=0$,
$\omega^{2}_{+}(0)$ $=$ $m^{2}_{>}\,,$
$\omega^{2}_{-}(0)$ $=$ $m^{2}_{<}\,,$ (7.1)
where $m_{>}$ ($m_{<}$) is the greater (lesser) of $m_{\rho}$ and $m_{V}$. In what follows we consider the case $m_{\rho}>m_{V}$. The opposite case can be treated similarly.
@@
A convenient way to write the elements of the mode eigenvectors in this case is
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}$ $=$ $\frac{1}{\sqrt{N_{-}}}\left(\begin{array}[]{c}m_{V}\gamma_{-}\\[12.0pt] \Lambda+\gamma^{2}_{-}\end{array}\right)\,,$ (7.6)
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{+}$ $=$ $\frac{1}{\sqrt{N_{+}}}\left(\begin{array}[]{c}\Lambda\\[12.0pt] -m_{V}\gamma_{+}\end{array}\right)\,,$ (7.11)
@@
$k^{2}_{\pm}$ $=$ $\omega^{2}_{\pm}-\kappa^{2}=\frac{1}{2}(m^{2}_{\rho}+m^{2}_{V})\pm\left[\frac{1}{4}(m^{2}_{\rho}-m^{2}_{V})^{2}+4\mu^{2}\kappa^{2}\right]^{\frac{1}{2}}\,,$
$\Lambda$ $=$ $\frac{1}{2}(m^{2}_{\rho}-m^{2}_{V})+\left[\frac{1}{4}(m^{2}_{\rho}-m^{2}_{V})^{2}+4\mu^{2}\kappa^{2}\right]^{\frac{1}{2}}\,,$ (7.12)
@@
$N_{-}$ $=$ $(\Lambda+\gamma^{2}_{-})^{2}+m^{2}_{V}|\gamma_{-}|^{2}\,,$
$N_{+}$ $=$ $\Lambda^{2}+m^{2}_{V}\gamma^{2}_{+}\,.$ (7.13)
With this convention, in the limit $\kappa=0$ these solutions reduce to
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}$ $=$ $\left(\begin{array}[]{c}0\\ 1\end{array}\right)\,,$ (7.18)
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{+}$ $=$ $\left(\begin{array}[]{c}1\\ 0\end{array}\right)\,,$ (7.23)
when $m_{\rho}>m_{V}$. Plots of the mixing coefficients $|\rho_{\pm}|^{2}$ and $|V_{L\pm}|^{2}$ is shown in Fig. 6 and Fig. 7. For the plots we take $M^{2}$ and $\Delta$ as indicated in the caption, and we consider the case $m_{\rho}>m_{V}$, so that, in particular, $(m_{V}/2\mu)^{2}=\frac{1}{2}(M^{2}-\Delta)$. The plots do not vary significantly for other values of $M^{2}$ and $\Delta$.
As already indicated above, at $\kappa=0$ the $\omega_{-}$ solution corresponds to a pure $V_{L}$,
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}=\left(\begin{array}[]{c}0\\ 1\end{array}\right)\qquad(\kappa=0)\,.$ (7.24)
At very large values of $\kappa$, it follows from Eq. (7.6) that the $\rho$ component dominates in both branches, which is reflected in the plots in Figs. 6 and 7.
@@
$\kappa_{0}=\frac{m_{V}m_{\rho}}{2\mu}\,.$ (7.26)
It is easy to see from Eq. (7), that $\gamma_{-}$ increases relative to $\Lambda$ as $k^{2}_{-}\rightarrow 0$, and therefore from Eq. (7.6) it follows that for $\kappa=\kappa_{0}$ the ()(-)( - ) mode is a pure $V_{L}$
@@
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}=\frac{1}{\sqrt{N_{-}}}\left(\begin{array}[]{c}-im_{V}|\gamma_{-}|\\[12.0pt] \Lambda-|\gamma_{-}|^{2}\,.\end{array}\right)\qquad(\kappa>\kappa_{0})\,.$ (7.29)
@@
$e_{\pm}=\left(\begin{array}[]{c}\pm ia\\ b\end{array}\right)\,.$ (7.30)
@@ 8 Conclusions and Outlook
We have presented a model for the propagation of photons in a medium that contains a Bose-Einstein (BE) condensate of a charged scalar field. The idea is based on the model recently proposed to treat the propagation of fermions in a BE condensate, extended by coupling the scalar field with the electromagnetic field in the usual way. Our goal has been to study the effects of the symmetry breaking on the photon dispersion relations and the optical properties of the system. To this end, the propagating modes were identified, and the corresponding dispersion relations were determined, in Section 4. As we have shown there, two of the modes correspond to the transverse photon polarizations, and their dispersion relations have the usual form of the transverse photons in a plasma. There are two other modes, to which we refer as the (±)plus-or-minus($\pm$)( ± ) modes, that are combinations of the longitudinal photon and the massive scalar field. Some salient aspects of the results obtained for the dispersion relations of the (±)plus-or-minus($\pm$)( ± ) modes were discussed in Section 5. The dispersion relation of the (+)(+)( + ) mode is always greater than the momentum, and increases steadily as the momentum increases. The behavior of the dispersion relation for the ()(-)( - ) mode is very different. On one hand, it decreases as the momentum increases in a given momentum range, and therefore the corresponding group velocity is negative in that range. In another range, for values of momentum above a given threshold, it increases steadily, but such that it remains smaller than the momentum as it increases, akin to a medium with index of refraction greater than 1. In Section 5 we also considered the non-relativistic limit of the dispersion relations, and discussed some aspects of the results in particular for the ()(-)( - ) mode. As shown there, under some conditions the dispersion relation of the ()(-)( - ) mode reduces to the form of the dispersion relation that has been obtained in non-relativistic models of the BE condensation of charged scalars[21, 22]. As a complement, in Section 6 we gave the expression for the contribution to the photon polarization tensor due to the BE condensate. This provides an alternative way to obtain the dispersion relations, but more importantly it has the virtue that it gives the corresponding contribution to the dielectric function of the system, which is useful on its own merits, e.g. for studying also the response of the system to external fields. Another useful result is the determination of the mixing parameters of the (±)plus-or-minus($\pm$)( ± ) modes presented in Section 7, which can be used to obtain the corrections to the dispersion relations (e.g., imaginary parts due the damping effects) and/or the effects of scattering, due to the interactions with the excitations of the system.
In summary, altogether, the results we have obtained show that the BE condensate model we have considered has some particular non-trivial optical properties. Our aim has been to analyze them, in the context of the model, with the expectation that the model and the results obtained can be useful in some of the specific application contexts considered in the references cited.
@@
What is the Lagrangian density in momentum space? Using the transformation of $a$ and $b$,
@@
In particular, notice that this $L(k)$ is not real (Hermitian). The correct procedure is to notice that, since the first term mixes $a(k)$ with $a(-k)$, and similarly with $b(k)$, we must include both also in the $\alpha$ term. Therefore we must write